The annual demand for a certain item is 22,500 pc/yr. One unit of the productcost $35.00, and the holding cost rate is 15%/yr. Setup time to produce a batch is 3.25 hr. The cost of equipment downtime during setup plus associate labor is $200/hr. Determine the economic order quantity and the total inventory cost for this case.
Answer #1

Given Information,

[ D ] Annual Demand = 22,500 PC/Yr,

One Unit of Product Cost = USD 35.00

[ H ] Holding cost Rate = 15% Per Year = 0.15 Per Year,

Setup time to produce a batch = 3.25 Hours

Cost of Equipment downtime during setup plus associate labor = USD 200/Hour,

So, [S] Cost per batch = USD 200 X 3.25 = USD 650 per batch

Answer:

Economic Order Quantity:

Formula is = Square Root of [( 2 X D X S)/H]

So, Economic Order Quantity: Square Root of [( 2 X 22,500 X 650)/0.15]

Economic Order Quantity = 13964 PC

Number of Batch or Order per year = Annual Demand / Economic Order Quantity

Number of Batch or Order per year = 22,500/13964 = 1.61 = 2

As Batch can’t be in the form of fraction Number,

Again,

Batch or Order Cost = Number of Batch or Order Per Year X Cost per Batch

Batch or Order Cost = 2 X USD 650 = USD 1,300

Holding Cost = Average Units X Holding cost Per Unit

Holding Cost = (1394/2) X 0.15 = USD 104.55

Total Product cost at Economic Order Quantity = USD 35.00 X 13964 = USD 488,740

Total Inventory Cost = ( Total Product cost at Economic Order Quantity + Holding Cost + Batch or Order Cost)

Total Inventory Cost = USD 488,740 + USD 104.55 + USD 1,300 = USD 490,144.55 = USD 490,145 (Answer)